OpenStudy (anonymous):

sin 645, what i am doing wrong

OpenStudy (anonymous):

645 means sine is negative (lies in third quadrant) sin(720-75) = -sin75 sin 75 = \[\frac{ \sqrt{3}-2 }{ 4\sqrt{2} }\]

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

I don't think that's sin(75) :)

OpenStudy (kainui):

I suggest subtracting 360 degrees, then reduce it by identities down to an angle between 0 and 90 degrees with the sign.

OpenStudy (aum):

-sin(75) = -sin(45 + 30)

OpenStudy (anonymous):

I mean... come on, sin(75) > 0 but... \[\sqrt3 - 2 < 0\]

OpenStudy (anonymous):

sin(45+30) sin45cos30 - cos45sin30

OpenStudy (anonymous):

Silly nameless...person... \[\Large \sin(A+B) = \sin(A)\cos(B) \color{red}+\cos(A)\sin(B)\]

OpenStudy (aum):

sin45cos30 + cos45sin30

OpenStudy (kainui):

Are you trying to evaluate sin(75) or are you trying to show that identity is true? I'm sort of confused now.

OpenStudy (anonymous):

oh yes , @kainui nothing i made a terrible mistake , i added a - sign

OpenStudy (anonymous):

maybe you're good to go now? ^^

OpenStudy (anonymous):

and not that I'm completely doubting your arithmetic skill, but... \[\large \sin(45) = \frac{\sqrt2}{2}=\cos(45)\]\[\large \sin(30) =\frac12 \]\[\large \cos(30) = \frac{\sqrt3}{2}\]

OpenStudy (anonymous):

thanks ! hehe

OpenStudy (anonymous):

My book has got sin 45 and cos as 1/root 2

OpenStudy (anonymous):

It's wrong

OpenStudy (aum):

It is the same. 1/sqrt(2) = sqrt(2) / 2

OpenStudy (aum):

some books and teachers don't like radical in the denominator.

OpenStudy (anonymous):

Silly :>

OpenStudy (anonymous):

i have not slept from a long time

OpenStudy (anonymous):

Then get some rest haha